# Stellar intensity interferometry in the photon-counting regime

### v1

@@ Abstract
I Short review on stellar intensity interferometry
I.1 Historical work by Hanbury Brown and Twiss
The history of intensity interferometry and photon correlation is closely related to the development of modern quantum optics. After Einstein introduced the concept of light quanta, the way to reconcile this corpuscular picture with interference phenomena was to consider that photons interfere with themselves. Until the 1950s, physicists widely accepted Dirac’s assertion that “interference between two different photons can never occur” [1]. However, this view was deeply transformed thanks to the pioneering work of Hanbury Brown and Twiss (HBT), whose experiments challenged traditional ideas. Their results on intensity correlations were finally recognized as key milestones towards our understanding of the quantum nature of light, and triggered the development of the quantum theory of optical coherence [2, 3, 4].
@@ I.1 Historical work by Hanbury Brown and Twiss
[…]
@@
where $V(r)$ is the “visibility”, i.e., the contrast of the interference fringes if the light from the two telescopes were combined. The first-order correlation is directly given by the Fourier transform of the light power spectrum (Wiener-Kinchin theorem), while the visibility is the Fourier transform of the luminosity distribution of the source (Zernike-van Cittert theorem). For example, for a uniform disk distribution of angular diameter $\theta$,
$V^{2}(r)=\left[\frac{2J_{1}(\pi r\theta/\lambda)}{\pi r\theta/\lambda}\right]^{2}\,,$ (4)
@@
However, these first results obtained by HBT were met with great skepticism. A notable feature of the early HBT experiments was that they were performed with continuous wave detection. In this regime, photons were not required to explain the observed results, as the classical wave theory of light is sufficient. The instantaneous intensity $I(t)$ could be interpreted as the probability of detecting light at a given moment, without invoking the concept of discrete photons. Nevertheless, when intensity interferometry was considered in the photon-counting regime, the classical framework could no longer fully account for the results, and a quantum description became necessary.
In 1956, Purcell provided a theoretical insight: photon bunching, observed in intensity correlation measurements with $g^{(2)}(\tau=0)>1$, was a natural consequence of the Bose-Einstein statistics obeyed by photons as bosons [9]. HBT further developed this interpretation [10], as did Kahn [11], and the effect was experimentally confirmed in a photon-counting experiment by HBT themselves [12]. Slightly later, Fano introduced another interpretation, framing the phenomenon in terms of two-photon interference [13]. Together, these contributions deepened the understanding of photon statistics and light’s quantum nature, and intensity correlation is now a tool that is widely used to probe the quantum nature of light.
The use of spatial intensity correlations was then widely applied by HBT in the 1960s and 1970s. Since the amplitude of the bunching peak $g^{(2)}(r,\tau=0)$, obtained by correlating the light detected by two telescopes separated by a projected baseline $r$, is directly proportional to the squared modulus of the visibility (Eqs. 2 and 3), one has access to the star’s diameter. This approach allowed the measurement of the angular diameter of 32 stars using the Narrabri Stellar Intensity Interferometer in Australia [14].
Despite its success, the technique was ultimately replaced by amplitude interferometry, pioneered by A. Labeyrie [15]. Unlike stellar amplitude interferometry, which involves directly combining light from multiple telescopes to produce interference fringes and demands precise control of path-length differences at the scale of an optical wavelength, stellar intensity interferometry (SII) is relatively phase-insensitive and simpler to implement. However, this simplicity comes at the cost of significantly lower sensitivity (i.e. signal-to-noise ratio) compared to amplitude interferometry, which has, to date, restricted its application to the observation of very bright stars [14].
I.2 Modern revival of intensity interferometry
@@
These arrays offer two key advantages. First, their large collectors, with diameters of more than 10 meters, enable the capture of a significant number of photons, improving the signal-to-noise ratio and allowing, in particular, access to fainter stars. Secondly, their spatial arrangement enables correlation measurements across a wide range of baselines with a two-dimensional distribution. While intensity interferometers can only measure the squared modulus of the electric field’s correlation function — thereby losing phase information and making direct star image reconstruction impossible  [^1] Accessing the Fourier phase is in principle feasible by measuring third order correlations [72, 73], but this has not yet been demonstrated due to a very low signal-to-noise ratio. — the use of an interferometer with varying baselines in two-dimensional space can impose significant constraints on the source image [16].
@@
including our group in Nice (see next section), have adopted this approach.
@@
Lastly, while the ultimate goal of the use of optical telescopes is to reach faint magnitudes (up to magnitude 8 for current experiments, higher in the future), sub-milliarcesond resolution, requiring baselines of up to a few kilometers (see, in particular, the project of a Swiss consortium [30]), a complementary strategy involves the use of large arrays of small telescopes [31]. Although the smaller aperture limits the signal-to-noise ratio and restricts observations to very bright stars, the ability to construct an array with a large number of telescopes and numerous short baselines could enable precise image reconstruction of bright giant stars [32, 33, 34]. Moreover, a dedicated array could be used for long observing programs to reach fainter magnitudes, as the signal from individual observations can be combined post-facto. Let us finally note that this idea of using many small apertures has also been proposed for Cherenkov telescopes by exploiting individual facets as independent sub-apertures and applying aperture synthesis techniques using pairwise and triple intensity correlations [35].
I.3 Results of the Nice group
Figure 2: Intensity interferometry performed by the I2C consortium at Calern Observatory (France). Left: The two telescopes of C2PU used for spatial intensity interferometry. Right: The coupling device connected to the telescope’s Cassegrain port. It spectrally filters the incoming light, separate the two polarization channels, and inject the light into multimode fibers that feed photon detectors [36].
Our group, a collaboration of physicists and astronomers within the Université Côte d’Azur network, began its work on SII in 2017, building on earlier research using intensity correlations to study light scattered by hot atomic vapors [37]. Since then, we have achieved several notable milestones. Thanks to the available facilities at Calern Observatory, we first observed the bunching peak in stellar light using a single telescope [38] and later performed SII with the two C2PU telescopes shown on the left of Fig. 2 [39]. These observations marked the first successful intensity interferometry measurements of stars  [^2] With a single telescope, photon bunching was observed with Sun light slightly earlier [52]. since the Narrabri Observatory era [14] and were the first to use photon-counting detection.
@@ I.3 Results of the Nice group
One long-term goal is to develop an instrument capable of performing interferometric measurements across visible wavelengths (B, V, R, I bands) with sensitivity to stars brighter than magnitude 8 when deployed on major facilities. Achieving this goal requires a significant enhancement of the sensitivity of the current instrument, for instance, through wavelength multiplexing and the use of even better detectors (see Section II.4). A critical aspect in evaluating the future potential is the determination of the signal-to-noise ratio.
II The signal-to-noise ratio with single-photon detectors
After this overview of the state-of-the-art, in this second part, we discuss in detail the signal-to-noise ratio (SNR). We first recall HBT’s well-known formula and its validity conditions, and then how to adapt it and derive it in the photon-counting regime. We show numerical and experimental data that validate our SNR equation. We also discuss experimental artifacts to be avoided to reach the theoretical SNR and some limitations that are hard to avoid. Finally, we discuss the on-going work to significantly enhance the sensitivity by implementing wavelength multiplexing with high-time resolution detectors.
II.1 HBT’s formula
@@
  • $N_{\mathrm{ph}}(\lambda)$ the photonic spectral flux from the star at wavelength $\lambda$, i.e. a number of photons per second, per squared meter and per Hertz of optical bandwidth (unit: ph.s-1.m-2.Hz-1).
  • $A$ the collecting area of each telescope, supposed to be equal. If they are not equal, the geometric mean should be used.
@@
  • The incident light is unpolarized.
  • Spurious light and dark current (dark counts for single-photon detectors) are negligible.
  • Spurious correlations are negligible.
  • The electronics act as a low-pass filter of infinite order and do not introduce any other noise (i.e., the only noise is the shot noise).
@@
This last assumption explains why the SNR is independent of the optical bandwidth $\Delta\nu$. Increasing $\Delta\nu$ also increases the number of detected photons but decreases the temporal coherence $\tau_{\mathrm{c}}=1/\Delta\nu$. Since the $g^{(2)}(\tau)$ function is convolved by the much larger electronic response, this reduces the contrast of the bunching peak. The two effects compensate exactly. In practice, spectral filtering is necessary to fulfill the above assumptions. Indeed, a too large spectrum may reduce the coherence time so much that the amplitude of the convolved bunching peak may become comparable to spurious correlations, and also the detectors may saturate. Moreover, the visibility depends on the wavelength (Eq.,4), so a broad spectrum makes the interpretation of the results more complex. On the other hand, too narrow filtering may reduce the photon rate from the source so much that the dark count or spurious light may become nonnegligible. Typically, spectral filtering between one and a few nm is used. Another remark is that polarizing the light also does not change the final SNR: it increases the bunching peak by a factor two, but decreases $\alpha$ by the same factor, corresponding to the transmission through the polarizer  [^3] Processing independently the two polarization channels leads to a $\sqrt{2}$ increases of the SNR at the expense of doubling the number of detectors [36, 45, 42].. Finally, this equation is adapted to the case of measuring correlation between two telescopes, each of collecting area $A$. If, for calibration purpose, one measures the so-called zero-baseline correlation, i.e. the $g^{(2)}(\tau,r=0)$ at a single telescope by splitting the light onto two detectors (like in our first experiment [38]), then the SNR of that measurement is divided by two.
@@
In the more general case of two telescopes and fluctuating count rates (due to seeing and scintillation), the SNR will be proportional to $<\sqrt{F_{1}(t)F_{2}(t)}>$ (average over time), which can be slightly smaller than $\sqrt{<F_{1}(t)><F_{2}(t)>}$ in case of strong uncorrelated fluctuations.
II.2 Adaptation to the photon-counting regime
In recent years, this formula has been often used by replacing $\Delta f$ with an “electronic timing resolution” $\tau_{\mathrm{el}}$, a terminology more appropriate in the photon-counting regime. Most of the time, it has been said that $\Delta f\sim 1/\tau_{\mathrm{el}}$, without precisely defining $\tau_{\mathrm{el}}$ (rms, FWHM) [38, 48] or with a precise definition, e.g. the time bin, but without questioning the numerical prefactor [27]. In this section, our goal is to clarify this point.
@@ II.2 Adaptation to the photon-counting regime
with $\tau_{\mathrm{el}}$ the rms width of the timing resolution of the correlation (i.e., $\sqrt{2}$ times the timing resolution of a single detector), assuming a Gaussian jitter distribution function for the detectors. Compared to HBT’s formula  [^5] It is easy to compute the Bode diagram corresponding to a Gaussian jitter, and one finds that the -3 dB bandwidth is $\Delta f=\sqrt{\ln(2)}/(2\pi\tau_{\mathrm{el}})$, a very slightly lower value than $\Delta f=1/(4\sqrt{\pi}\tau_{\mathrm{el}})$. This is because Eq. (5) is simplified (only valid for an infinite-order filter, see. [46])., the electronic bandwidth $\Delta f$ has been replaced by $(4\sqrt{\pi}\tau_{\mathrm{el}})^{-1}$. We have highlighted the factor $1/2$ due to the fact that light is unpolarized, it should be removed if light is polarized.
@@
II.2.1 Alternative derivations of the SNR
Let us consider an HBT setup with two detectors in the photon-counting regime, each connected to a time-to-digital convertor (TDC, or time-tagger) that records the arrival time of each detected photon (see Fig. 4(a) for an illustration). The unnormalized $G^{(2)}(\tau)$ is the histogram of all delays $\tau$ between counts on the two detectors. In the absence of any correlation (i.e. far from the bunching peak), on average, the accumulated number of coincidences in each time bin of the histogram is

### v2

@@ Abstract
keywords:
Stellar interferometry, Hanbury Brown and Twiss Effect, Single-photon detectors
{altabstract}
L’interférométrie d’intensité stellaire consiste à mesurer la corrélation entre les fluctuations d’intensité de la lumière captée par deux télescopes observant la même étoile. L’amplitude de la corrélation est directement liée à la taille de l’étoile, qui ne serait pas résolue par un seul télescope. Cette technique est basée sur le célèbre effet Hanbury Brown et Twiss. Après sa découverte dans les années 1950, elle a été utilisée en astronomie jusque dans les années 1970, puis remplacée par l’interférométrie directe (ou d’amplitude), qui est beaucoup plus sensible, mais aussi beaucoup plus exigeante techniquement. Cependant, depuis quelques années, l’interférométrie d’intensité connait un renouveau. Dans cet article, nous présentons un résumé de l’état de l’art et nous discutons en détail du rapport signal à bruit de l’interférométrie d’intensité dans le cadre de détecteurs en comptage de photons.
1 Short review on stellar intensity interferometry
1.1 Historical work by Hanbury Brown and Twiss
The history of intensity interferometry and photon correlation is closely related to the development of modern quantum optics. After Einstein introduced the concept of light quanta, the way to reconcile this corpuscular picture with interference phenomena was to consider that photons always interfere with themselves. Until the 1950s, the notion of multiparticle interference was unknown to most physicists. This view was deeply transformed thanks to the pioneering work of Hanbury Brown and Twiss (HBT)[1, 2], whose experiments challenged traditional ideas. Their results on intensity correlations were finally recognized as key milestones towards our understanding of the quantum nature of light, and triggered the development of the quantum theory of optical coherence [3, 4, 5].
@@ 1.1 Historical work by Hanbury Brown and Twiss
with $I_{1}$ and $I_{2}$ the intensity collected by the two telescopes, $\tau$ the time lag between them, $r$ the projected baseline between the telescopes, and with $\langle\,\cdot\,\rangle$ corresponding to the averaging over time $t$. For chaotic light, a classical explanation can relate this intensity correlation function to the electric field correlation function,
$g^{(1)}(\tau,r)=\frac{\langle E_{1}^{\star}(t)E_{2}(r,t+\tau)\rangle}{\langle\sqrt{I_{1}I_{2}}\rangle}\,.$ (2)
The so-called Siegert relation [6, 7] states :
@@
where $V(r)$ is the “visibility”, i.e., the contrast of the interference fringes if the light from the two telescopes were combined. The first-order time correlation is directly given by the Fourier transform of the light power spectrum (Wiener-Khinchin theorem), while the visibility as a function of the spatial separation $r$ is the Fourier transform of the luminosity distribution of the source (van Cittert-Zernike theorem) [8]. For example, for a uniform disk distribution of angular diameter $\theta$,
$|V(r)|^{2}=\left[\frac{2J_{1}(\pi r\theta/\lambda)}{\pi r\theta/\lambda}\right]^{2}\,$, (5)
@@
However, these first results obtained by HBT were met with great skepticism (see [9] for a historical review of the controversy). A notable feature of the early HBT experiments was that they were performed with continuous wave detection. In this regime, photons were not required to explain the observed results, as the classical wave theory of light is sufficient. The instantaneous intensity $I(t)$ could be interpreted as the probability of detecting light at a given moment, without invoking the concept of discrete photons. Nevertheless, when intensity interferometry was considered in the photon-counting regime, the classical framework could no longer fully account for the results, and a quantum description became necessary.
In 1956, Purcell provided a theoretical insight : photon bunching, observed in intensity correlation measurements with $g^{(2)}(r=0,\,\tau=0)>1$, was a natural consequence of the Bose-Einstein statistics obeyed by photons as bosons [10]. HBT further developed this interpretation [11], as did Kahn [12], and the effect was experimentally confirmed in a photon-counting experiment by HBT themselves [13] and by Pound & Rebka [14]. Slightly later, Fano introduced another interpretation, framing the phenomenon in terms of two-photon interference [15]. Together, these contributions deepened the understanding of photon statistics and lightâs quantum nature, and intensity correlation is now a tool that is widely used to probe the quantum nature of light.
The use of spatial intensity correlations was then widely applied by HBT in the 1960s and 1970s. Since the amplitude of the bunching peak $g^{(2)}(r,\tau=0)$, obtained by correlating the light detected by two telescopes separated by a projected baseline $r$, is directly proportional to the squared modulus of the visibility (Eqs. 3 and 4), one has access to the star’s diameter. This approach allowed the measurement of the angular diameter of 32 stars using the Narrabri Stellar Intensity Interferometer in Australia [16].
Despite its success, the technique was ultimately replaced by amplitude interferometry, pioneered by A. Labeyrie, who obtained first fringes on Vega in 1975 at Calern (France) [17]. Unlike stellar amplitude interferometry, which involves directly combining light from multiple telescopes to produce interference fringes and demands precise control of path-length differences at the scale of an optical wavelength, stellar intensity interferometry (SII) is phase-insensitive and simpler to implement. However, this simplicity comes at the cost of significantly lower sensitivity (i.e. signal-to-noise ratio) compared to amplitude interferometry, which has, to date, restricted its application to the observation of very bright stars [16].
1.2 Modern revival of intensity interferometry
@@
These arrays offer two key advantages. First, their large collectors, with diameters of more than 10 meters, enable the capture of a significant number of photons, improving the signal-to-noise ratio and allowing, in particular, access to fainter stars. Secondly, their spatial arrangement enables correlation measurements across a wide range of baselines with a two-dimensional distribution. While intensity interferometers can only measure the squared modulus of the electric field’s correlation function â thereby losing phase information and making direct star image reconstruction impossible [^1] Accessing the Fourier phase is in principle feasible by measuring third order correlations [19, 20], but this has not yet been demonstrated due to a very low signal-to-noise ratio. â the use of an interferometer with varying baselines in two-dimensional space can impose significant constraints on the source image [18].
@@
Lastly, while one of the prime goals of the use of optical telescopes is to reach faint magnitudes (up to magnitude 8 for current experiments, higher in the future) and sub-milliarcesond resolution, requiring baselines of up to a few kilometers (see, in particular, the project of a Swiss consortium [33]), a complementary strategy involves the use of large arrays of small telescopes [34]. Although the smaller aperture limits the signal-to-noise ratio and restricts observations to very bright stars, the ability to construct an array with a large number of telescopes and numerous short baselines could enable precise image reconstruction of bright giant stars [35, 36, 37]. Moreover, a dedicated array could be used for long observing programs to reach fainter magnitudes, as the signal from individual observations can be combined post-facto. Let us finally note that this idea of using many small apertures has also been proposed for Cherenkov telescopes by exploiting individual facets as independent sub-apertures and applying aperture synthesis techniques using pairwise and triple intensity correlations [38].
1.3 Results of the Nice group
Our group, a collaboration of physicists and astronomers within the Université Côte dâAzur network, began its work on SII in 2017, building on earlier research using intensity correlations to study light scattered by hot atomic vapors [39]. Since then, we have achieved several notable milestones. Thanks to the available facilities at Calern Observatory, we first observed the bunching peak in stellar light using a single telescope [40] and later performed SII with the two C2PU telescopes (15 m baseline) shown on the left of Fig. 2 [41]. These observations marked the first successful intensity interferometry measurements of stars [^2] With a single telescope, photon bunching was observed with Sun light slightly earlier by Tan and coworkers [42]. since the Narrabri Observatory era [16] and were the first to use photon-counting detection.
@@ 1.3 Results of the Nice group
Figure 2: Intensity interferometry performed by the I2C consortium at Calern Observatory (France). Left : The two telescopes of C2PU, separated by 15 m, used for spatial intensity interferometry. Right : The coupling device connected to the telescopeâs Cassegrain port. It spectrally filters the incoming light, separates the two polarization channels, and injects the light into multimode fibers that feed photon detectors [44].
One long-term goal is to develop an instrument capable of performing interferometric measurements across visible wavelengths (B, V, R, I bands) with sensitivity to stars brighter than magnitude 8 when deployed on major facilities. Achieving this goal requires a significant enhancement of the sensitivity of the current instrument, for instance, through wavelength multiplexing and the use of even better detectors (see Section 2.4). A critical aspect in evaluating the future potential is the determination of the signal-to-noise ratio.
2 The signal-to-noise ratio with single-photon detectors
After this overview of the state of the art, in this second part, we discuss in detail the signal-to-noise ratio (SNR). We first recall HBT’s well-known formula and its validity conditions, and then how to adapt it and derive it in the photon-counting regime. We show numerical and experimental data that validate our SNR equation. We also discuss experimental artifacts to be avoided to reach the theoretical SNR and some limitations that are hard to avoid. Finally, we discuss the on-going work to significantly enhance the sensitivity by implementing wavelength multiplexing with high-time resolution detectors.
2.1 HBT’s formula
@@
  • $N_{\mathrm{ph}}(\lambda)$ the photonic spectral flux from the star at wavelength $\lambda$, i.e. a number of photons per second, per squared meter and per Hertz of optical bandwidth (unit : ph s-1 m-2 Hz-1).
  • AA the collecting area of each telescope, supposed to be equal. If they are not equal, the geometric mean should be used.
@@
  • the incident light is unpolarized ;
  • spurious light and dark current (dark counts for single-photon detectors) are negligible ;
  • spurious correlations are negligible ;
  • the electronics act as a low-pass filter of infinite order and do not introduce any other noise (i.e., the only noise is the shot noise) ;
@@
This last assumption explains why the SNR is independent of the optical bandwidth $\Delta\nu$. Increasing $\Delta\nu$ also increases the number of detected photons but decreases the temporal coherence $\tau_{\mathrm{c}}=1/\Delta\nu$. Since the measured $g^{(2)}(\tau)$ function is the convolution of the true $g^{(2)}(\tau)$ before instrumental effect, of width $\tau_{\mathrm{c}}$, with the electronic response of width $1/\Delta f$, the contrast of the bunching peak is reduced to $\sim\Delta f/\Delta\nu\ll 1$. The two effects compensate exactly. In practice, spectral filtering is necessary to fulfill the above assumptions. Indeed, a too large spectrum may reduce the coherence time so much that the amplitude of the convolved bunching peak may become comparable to spurious correlations, and also the detectors may saturate. Moreover, the visibility depends on the wavelength (Eq.,5), so a broad spectrum (large $\Delta\lambda$) makes the interpretation of the results more complex. On the other hand, too narrow filtering may reduce the photon rate from the source so much that the dark count or spurious light may become nonnegligible. Typically, spectral filtering between one and a few nm is used. Another remark is that polarizing the light also does not change the final SNR : it increases the bunching peak by a factor two, but decreases $\alpha$ by the same factor, corresponding to the transmission through the polarizer [^3] Processing independently the two polarization channels leads to a $\sqrt{2}$ increase of the SNR at the expense of doubling the number of detectors [44, 48, 45].. Finally, this equation is adapted to the case of measuring correlation between two telescopes, each of collecting area $A$. If, for calibration purpose, one measures the so-called zero-baseline correlation, i.e. the $g^{(2)}(\tau,r=0)$ at a single telescope by splitting the light onto two detectors (like in our first experiment [40]), then the SNR of that measurement is divided by two.
@@
In the more general case of two telescopes and fluctuating count rates (due to seeing and scintillation), the SNR will be proportional to $\sqrt{F_{1}(t)F_{2}(t)}>$ (average over time), which can be slightly smaller than $\sqrt{<F_{1}(t)><F_{2}(t)>}$ in case of strong uncorrelated fluctuations.
2.2 Adaptation to the photon-counting regime
In recent years, this formula has been often used by replacing $\Delta f$ with an “electronic timing resolution” $\tau_{\mathrm{el}}$, a terminology more appropriate in the photon-counting regime. Most of the time, it has been said that $\Delta f\sim 1/\tau_{\mathrm{el}}$, without precisely defining $\tau_{\mathrm{el}}$ (for example : rms or FWHM ?) [40, 50] or with a precise definition, e.g. the time bin, but without questioning the numerical prefactor [30]. In this section, our goal is to clarify this point.
@@ 2.2 Adaptation to the photon-counting regime
with $\tau_{\mathrm{el}}$ the rms width of the timing resolution of the correlation (i.e., $\sqrt{2}$ times the timing resolution of a single detector), assuming a Gaussian jitter distribution function for the detectors. Compared to HBT’s formula [^5] It is easy to compute the Bode plot corresponding to a Gaussian jitter, and one finds that the -3 dB bandwidth is $\Delta f=\sqrt{\ln(2)}/(2\pi\tau_{\mathrm{el}})$, a very slightly lower value than $\Delta f=1/(4\sqrt{\pi}\tau_{\mathrm{el}})$. This is because Eq. (6) is simplified (only valid for an infinite-order filter, see. [49])., the electronic bandwidth $\Delta f$ has been replaced by $(4\sqrt{\pi}\tau_{\mathrm{el}})^{-1}$. We have highlighted the factor $1/2$ due to the fact that light is unpolarized, it should be removed if light is polarized.
@@
2.2.1 Alternative derivations of the SNR
Let us consider an HBT setup with two detectors in the photon-counting regime, each connected to a time-to-digital convertor (TDC, or time tagger) that records the arrival time of each detected photon (see Fig. 4(a) for an illustration). The unnormalized $G^{(2)}(\tau)$ is the histogram of all delays $\tau$ between counts on the two detectors. In the absence of any correlation (i.e. far from the bunching peak), on average, the accumulated number of coincidences in each time bin of the histogram is